10. Find the center of gravity of a rod 50cm long, half of which is made of iron and half of which is made of aluminum.


1
Expert's answer
2021-03-29T09:04:35-0400
"x_G=\frac{0.25\rho _F+0.075\rho_A}{\rho _F+\rho_A }L\\
x_G=\frac{0.25(7874)+0.75(2712)}{(7874)+(2712)}50=18.9\ cm"